ACM EMAIL ROOM FAQ

WHAT SHOULD I DO IF THERE IS A PROBLEM? Please ask the staffer for help. Do not reboot the machines or do anything in the control panel.

WHERE IS THE MOUSE? These computers have a touchstick in the middle of the keyboard between the ’g’, ’h’, and ’b’ keys. This moves the mouse pointer. The mouse buttons are below the space bar. The left button is like the left button on a mouse. The right button is like the right button on a mouse. So, if you right-click on things, it’s different from left-clicking on things, just like with a mouse. If you click and get a menu you didn’t want, try left-clicking in a remote part of the desktop.

HOW DO I START NETSCAPE? Get rid of all other windows by clicking on the × button in their upper-right corners. Then double-click the Netscape Navigator shortcut to start Netscape.

WHAT SHOULD I DO WHEN I AM DONE? Please leave the computers running Netscape. When you are done with telnet, please kill its window by clicking on the × button in the title bar. The staff and the next user thank you!

DO YOU HAVE SSH? No. SSH is a secure remote login system that uses public-key encryption to protect your password from network sniffers, who break system security by watching network traffic and stealing passwords as they go by. It is an excellent system but it requires special software on both ends of the connection. To allow users to login from remote sites that do not have SSH installed, most sites also install a one-time password system, such as S/Key (http://www.gbnet.net/public/security/ nrl-opie/opie-2.32.tar.gz), which requires no special setup on the transmitting end. The remote user connects with ordinary telnet or rlogin and is challenged for a password from a list of 100 passwords the user carries. The user enters the requested password and is logged in. That password is never used again, so it is not useful to sniffers.

HOW DO I SEND EMAIL? Either connect to your home machine and send from there, or double-click on Bear Access and select Mail. This will start Eudora. Please quit Eudora when you are done and leave the machine running Netscape (see above).

WHY DOESN’T MY HOME MACHINE TRUST THE MACHINES HERE? The machines here are on temporary addresses that are not registered with the Domain Name Service (DNS). Some sites do not allow connections from such machines. DNS registration is not a requirement for connection to the Internet, so sites limiting access based on DNS registration risk locking their users out when they are travelling.

HOW DO I PRINT A POSTSCRIPT FILE ON A WEB SITE? Download the file by clicking on it in Netscape. A dialog box will appear. Click on ’Save File’. Save to the desktop. In the ’Start’ menu, select ’Programs’ and ’DOS Prompt’. A window will appear.

Type ’cd desktop’. Then enter ’type whatever.ps >lpt1:’ (replace “whatever.ps” with the name of your file). Click on the × button to close the DOS window and drag the file to the ’Recycle bin’ icon when you are done with it.

CAN YOU HELP ME GET THE SOFTWARE ON THIS FLOPPY TO WORK? No.

HOW DO I PRINT FROM MY PERSONAL LAPTOP?

Linux: As root, type ’printtool’, click on \underline{Add}, choose \underline{Remote Unix (lpd) Queue}, in the dialog box that appears, enter:

Names  acm
Spool Directory /var/spool/lpd/acm
File Limit in Kb 0
Remote Host  132.236.46.14
Remote Queue raw

Click on \underline{Select}, in the new dialog box, under ’Printer Type’, scroll down to ’PostScript printer’ and click on it. Click ’OK’. Click ’OK’ in the previous dialog box. You should be set up to print. In the ’PrintTool’ menu, choose ’Quit’.

PC or Mac: We’re working on it...